How many atoms of copper are present in a 4.8 g sample of copper metal?

Express your answer to two significant figures.

TIP: To report an answer in scientific notation, enter it using the format "2.3E4", which means "2.3 x 104" (without the quotation marks)
Answer #1

Given amount of Copper = 4.8 g

So mols of Copper present = mass / molar mass of Copper

= 4.8 g / 63.5 g/mol

= 0.075 mols

Again each mole contain = 6.023 * 1023 atoms /mole

So 0.075 mols Copper contain = 0.075 mols * 6.023 * 1023 atoms /mole

= 0.456 * 1023 atoms

= 4.56 * 1022 atoms

= 4.6 * E22 atoms